# The Prompt Makes the Person(a): A Systematic Evaluation of Sociodemographic Persona Prompting for Large Language Models

### v1

@@ 4 Results
We now report the results of our analyses of the five different LLMs by first focusing on (i) Demographic Representativeness, i.e., how well the LLMs simulate different identity groups for the three different tasks and their corresponding evaluation metrics. Then, we assess the impact of (ii) Strategies for Persona Prompting. Finally, we compare the five different models and explore whether some are more representative than others. For the open-ended tasks, we show results for the Self-Description task in the main body and report complementary results for the Social Media Bio task in Appendix E.2.
@@ 4.1 Demographic Representativeness
To assess how LLMs simulate different demographic groups, we first analyze the average number of identified marked words and the average semantic diversity of all demographic groups. Figure 2 shows that self-descriptions associated with nonbinary personas exhibit significantly lower semantic diversity and more marked words, while responses for male personas are among the most semantically diverse. We further observe that responses associated with Middle Eastern and Hispanic personas tend to contain more marked words and exhibit lower semantic diversity as opposed to other race/ethnicity groups. 666We see similar results for social media bios and report those results in Figure 8 in Appendix E.2. This suggests that the portrayal of marginalized gender and race/ethnicity groups are more prone to flattened portrayals, i.e., reduced diversity and recurring, possibly stereotypical patterns.
Next, we examine whether LLMs produce non-English responses during persona simulation. Across all LLMs, we find that non-English self-descriptions and bios are most frequent for Hispanic personas (up to ), while non-English responses are rare for all other racial groups (up to ). This indicates that LLMs fall victim to the ‘perpetual foreigner stereotype’ when simulating Hispanic personas (Dennis, 2018).
Finally, for OpinionsQA, we compare the average distributional alignment between simulated demographic personas and the human ground-truth (Fig. 3). In line with Suh et al. (2025), we observe the highest alignment for Black personas and the lowest alignment for White personas. However, we also find that the alignment achieved by two out of five models is worse than that of the random baseline (0.25 0.002; cf. Section 3.3.2). We discuss this observation in more detail in Section 4.3.
@@ 4.2 Strategies for Persona Prompting
We analyze prompt types using the Marked Words framework and find that prompts incorporating names for demographic priming and the interview format for role adoption produce the fewest marked words in self-descriptions (Fig. 4(a)) across all models. Beyond the Marked Words framework, we find that prompts using names and the interview format also promote overall higher semantic diversity (Fig. 4(b)).
Further, Figure 5 shows that prompts using names and the interview format lead to near-zero rates of non-English responses. Notably, the interview format seems to mitigate the language-switching effect introduced by explicit demographic priming. 777The reported patterns, specifically that names and the interview format lead to to improvements for the open-text measures, also hold for classification accuracy (see Fig. 12 in Appendix E.3) and for the Social Media Bio task, with detailed results in Figures 9 and 10 in Appendix E.2.
Finally, Figure 3 shows the distribution match in OpinionsQA (i.e., average Wasserstein distance) for the best performing model, OLMo-2-7B. We note a significant performance gain (i.e., lower Wasserstein distance) using the interview format.
Overall our findings indicate that using names for demographic priming and/or the interview format for role adoption leads to generally better representation of demographic groups across all models.888Further robustness checks and statistical significance for these results can be found in Appendix F.1.
To assess whether the observed positive effects benefit all groups equally, we compute the standard deviation of each measure across demographic groups. Table 1 shows that prompts using names and the interview format significantly reduce the standard deviation in marked word counts, indicating reduced disparities between demographic groups. This pattern generalizes across all open-text measures for OLMo-2-32B and OLMo-2-7B, indicating that these prompt types improve both overall representation and reduce inter-group disparities. Effects for other models are more mixed, suggesting that these prompt strategies sometimes benefit certain groups more than others. We report results for the remaining measures and all models in Table 8 in Appendix E.4.
@@ 4.3 Model Comparison
Surprisingly, two of the largest models, i.e., Llama-3.3-70B and Gemma-3-27B, perform worst with respect to the Marked Words framework and semantic diversity, while the smaller OLMo-2-7B, performs best (Fig. 4). 999We observe the same pattern for classification accuracy (see Fig. 12 in Appendix E.3) and for the Social Media Bio task (Fig. 9 in Appendix E.2).
@@ 5 Analysis of Marked Words
To complement our quantitative findings in Section 4, we also analyze the marked words identified in self-descriptions across all combinations of demographic group and prompt type. To this end, we compute the most frequent marked words per demographic group and present the top 10 in Table 2. In line with Cheng et al. (2023a), we observe that many of these words reflect patterns of markedness, essentialism, and othering in LLM responses. In particular, we identify four categories of marked words linked to problematic stereotypes as detailed in their work:
@@
Narrative of imposed resilience for Black personas (e.g. “resilience”, “strength”)
@@
Relationship to demographic identity as the primary frame of any non-White groups (e.g. “heritage”, “culture”)
@@
Conflation of Middle-Eastern identity with religiosity (e.g. “faith”, “muslim”)
@@
Disproportionate focus on gender identity for nonbinary personas (e.g. “gender”, “identity”)
For each category, we calculate the share of personas from the associated groups whose self-descriptions include these words and compare across prompt strategies. Figure 6 shows results for category 1 and 2: using names and/or the interview format consistently lowers the share of stereotyped persona descriptions across models.101010Results for additional stereotype categories follow the same pattern; full results and significance levels are provided in Appendix E.6.
@@ 6 Related Work
Personas in LLMs. Personas have been widely used in Interaction Design to represent user archetypes Cooper et al. (2014). They have also been adapted when molding the behavior of language technology, such as chatbots and more recently, LLMs. Personas in LLMs span a wide variety of dimensions, e.g., psychometric, occupational, or demographic and guide the model’s behavior and responses to better fit specific contexts or user needs Tseng et al. (2024). Inducing personas in LLMs effectively is an active area of research Jiang et al. (2023); Shu et al. (2024); Liu et al. (2024). We study demographic personas which can be used to create synthetic samples of a particular human subpopulation Argyle et al. (2023).
@@ 7 Discussion and Conclusion
Our findings show that LLMs often stereotype marginalized identities, especially nonbinary, Hispanic, and Middle Eastern identities, highlighting ongoing challenges in representing demographic groups (Cheng et al., 2023b, a).
@@
While name-based prompting also shows promise in mitigating biases, we advise caution in its use. In our analysis, we observe that models can misinterpret last names in unintended ways (for example, names associated with Black identity in the U.S. were sometimes interpreted as French), potentially compromising validity and skewing downstream analyses. Wang et al. (2025) used first names to signal gender identity, which can entrench biases towards certain groups, e.g, nonbinary people Gautam et al. (2024). We circumvent this to a great extent by relying on titles (i.e, Mr., Ms., and Mx.) to signal gender. Nonetheless, we suggest treating names as a potentially useful but ethically fraught tool, that requires careful implementation and critical reflection on its limitations.
Finally, we note that quantitative measures such as semantic diversity and marked word frequency should be considered along with qualitative analyses, task context, and normative goals. We encourage researchers to clearly document and justify their choice of role adoption, demographic priming, and specific descriptors used in prompts.
@@
While our work lays a foundation for evaluating sociodemographic persona prompts, several limitations remain. First, the study is limited to English and focuses only on two commonly studied dimensions — gender and race (Sen et al., 2025) — which constrains generalizability to other demographic dimensions and languages. Further, because our persona prompts are created in a fully data-driven manner, the resulting list is not exhaustive and may miss other meaningful variations. This limitation also applies to the choice of demographic descriptors; for example, we do not systematically compare terms like "girl" and "father", leaving such analysis for future work.
Name-based prompting presents additional challenges. Our racial categories and sets of last names are based on U.S. Census data, biasing the study toward U.S.-centric naming conventions and limiting applicability to other geographic and cultural contexts. Moreover, focusing only on the most common last names within each group risks capturing a narrow and potentially unrepresentative subset of the group’s identity. LLMs also do not always interpret names as intended. For instance, last names associated with Black identity in the U.S. were occasionally misinterpreted as French (e.g., “Mr. Jeanbaptiste, a 35-year-old history enthusiast from Paris”), indicating that name-based priming may invoke unintended geographic or cultural associations. Crucially, we caution against relying on first names to infer or encode gender, as this risks misgendering individuals and excluding nonbinary and gender-diverse people (Gautam et al., 2024); therefore, we use titles instead to signal gender.
We further do not control for the order of answer options in subjective question answering, due to computational constraints. This can influence response patterns and may impact the observed alignment behavior. Regarding potential data contamination in OpinionsQA, we considered this possibility but found it unlikely to be a confounding factor, as none of the models demonstrate particularly strong alignment on the task.
Further, we always place our persona prompts in the user message, as our literature review found this to be the most common approach in prior studies. We anticipate that using the system prompt instead could reduce the frequency of role violations, particularly for Gemma-3-27b, which showed the highest rate of such violations (cf. Appendix E.1). A systematic comparison of both strategies is an interesting venue for future work.
@@ Acknowledgments
We thank Florian Lemmerich for the valuable input and helpful discussions in the early stages of this project.
@@ Appendix B Sociodemographic Prompt Construction
Table 3 presents all prompt templates used in our experiments. Task-specific strings are indicated by a placeholder ([TASK]) and substituted as shown in Table 4. Placeholders indicating race/ethnicity through explicit descriptors ([R]) or last names ([NM]) are listed in Table 5. We source last names for all racial and ethnic groups from Aher et al. (2023), with the exception of Middle Eastern names, which are not included in that paper. Instead, we sample the most common surnames from the three largest Middle Eastern countries, namely Turkey,111111https://forebears.io/turkey/surnames Iran,121212https://forebears.io/iran/surnames and Saudi-Arabia.131313https://forebears.io/saudi-arabia/surnames We use all 10 names for each group in every task. For the open-ended tasks, we repeat the prompt with each individual name for 10 random seeds (instead of 100), resulting in a total of 100 responses per name category. The remaining placeholders are populated with demographic descriptors as indicated in Tables 6 and 7.
@@ Appendix C Computational Details
We use vLLM141414https://github.com/vllm-project/vllm version 0 (for the open-ended tasks) and version 1 (for OpinionsQA since it was faster) for LLM inference. We run the Llama-3.1-8B-Instruct and OLMo-2-1124-7B-Instruct models on a single Nvidia A100 (40GB) GPU with a total running time of approximately 9:50h (0:30h for self-descriptions, 0:20h for social media bios and 9:00h for OpinionsQA) per model. We run Llama-3.3-70B-Instruct, OLMo-2-0325-32B-Instruct and gemma-3-27b-it on two Nvidia H100 (80GB) GPUs in parallel. The total running time for generating all outputs with all three models is approximately 43:40h (5:40h for self-descriptions, 2h for social media bios, 36h for OpinionsQA).
For all models, we use the parameters recommended by the model creators. To generate answer option distributions for OpinionQA, we extract temperature-scaled log probabilities, applying the temperature values recommended by the model creators.
@@ Appendix D Preprocessing of LLM Responses
For the free-text responses we apply a variety of measures to identify stereotypes and flattening effects, with the normative expectation that groups are not treated differently when generating self-descriptions or social media bios for them based on demographics alone. To accurately detect such patterns, we first remove all demographic markers, names, and pronouns from our prompts that have been repeated in the LLMs’ answer. Using a regex, we replace them with the string "[identity]", while ensuring that consecutive redactions are combined. For instance, we replace "I am Ms. Garcia, a …" with "I am [identity], a …", when "Ms." and "Garcia" were both part of our prompt.
@@ D.1 Identification of Role Violations
We further filter the set of free-text responses to exclude cases in which the LLMs failed to maintain their assigned role. This includes instances where the model refused to answer (e.g. "I cannot fulfill this request."), answered as an AI assistant (e.g. "Just a language model navigating the human world.") or partially reverted to an AI identity after initially adopting the persona (e.g. "I am a 32-year-old Black female. I identify as an AI…").
To identify such role violations, we used Qwen2.5-32B Bai et al. (2023) - a model not included in our main experiments to avoid self-preference bias in LLM annotations Wataoka et al. (2024) - for automated annotation. We then validated this automatic classification by conducting a small annotation study, with 300 randomly sampled LLM responses where three human annotators who are fluent in English verified the labels. Both, the human annotators and Qwen2.5-32B received a similar instruction with the grey part being only shown to the LLM:

### v2

@@ 4 Results
We now report the results of our analyses of the five different LLMs by first focusing on (i) Demographic Representativeness, i.e., how well the LLMs simulate different identity groups for the two types of tasks and their corresponding evaluation metrics. Then, we assess the impact of (ii) Strategies for Persona Prompting. Finally, we compare the five different models and explore whether some are more representative than others. For the open-ended tasks, we show results for the Self-Description task in the main body and report complementary results for the Social Media Bio task in Appendix E.2.
@@ 4.1 Demographic Representativeness
To assess how LLMs simulate different demographic groups, we first analyze the average number of identified marked words and semantic diversity of demographic groups. Figure 2 shows that self-descriptions associated with nonbinary personas exhibit significantly lower semantic diversity and more marked words, while responses for male personas are among the most semantically diverse. We further observe that responses associated with Middle Eastern and Hispanic personas tend to contain more marked words and exhibit lower semantic diversity as opposed to other race/ethnicity groups. 777We see similar results for social media bios and report those results in Figure 8 in Appendix E.2. This suggests that simulations of marginalized gender and race/ethnicity groups are more prone to flattened portrayals, i.e., reduced diversity and recurring, possibly stereotypical patterns.
Next, we examine whether LLMs are prone to language switching during persona simulation. Across all LLMs, we find that non-English self-descriptions and bios are most frequent for Hispanic personas (up to ), while non-English responses are rare for all other racial groups (up to ). This indicates that LLMs fall victim to the ‘perpetual foreigner stereotype’ when simulating Hispanic personas (Dennis, 2018).
Finally, we compare the opinion distance between simulated personas and the human ground-truth on OpinionsQA (Fig. 3). In line with Suh et al. (2025), we observe the lowest distance for Black personas and the highest distance for White personas. However, we also find that two of the five models yield greater opinion distance than the random baseline (0.25 0.002). We discuss this in more detail in Section 4.3.
@@ 4.2 Strategies for Persona Prompting
We analyze prompt types using the Marked Personas framework and find that prompts incorporating names for demographic priming and the interview format for role adoption produce the fewest marked words in self-descriptions across all models (Fig. 4(a)). Beyond the Marked Personas framework, we find that prompts using names and the interview format also promote overall higher semantic diversity (Fig. 4(b)).
Figure 5 further shows that prompts using names and the interview format lead to near-zero rates of non-English responses. Notably, the interview format seems to mitigate the language-switching effect introduced by explicit demographic priming. 888The reported patterns, specifically that names and the interview format lead to to improvements for the open-text measures, also hold for classification accuracy (see Fig. 12 in Appendix E.3) and for the Social Media Bio task, with detailed results in Figures 9 and 10 in Appendix E.2.
Finally, Figure 3 shows the opinion distance in OpinionsQA (i.e., average Wasserstein distance) for the best performing model, OLMo-2-7B. We note a performance gain (i.e., lower Wasserstein distance) when using the interview format, indicating significantly reduced opinion distance.
Our findings indicate that using names for demographic priming and/or the interview format for role adoption leads to better representation of demographic groups across all models.999Further robustness checks and statistical significance for these results can be found in Appendix F.1.
To assess whether the observed positive effects benefit all groups equally, we compute the standard deviation of each measure across demographic groups. For each LLM, we run OLS regressions with this standard deviation as the dependent variable, analyzing the effect of prompt types. Table 1 reports the results for the marked word count, showing that prompts using names and the interview format significantly reduce the standard deviation across groups, indicating reduced disparities between demographic groups. This generalizes across all open-text measures for OLMo-2-32B and OLMo-2-7B, indicating that these prompt types improve both overall representation and reduce inter-group disparities. Effects for other models are more mixed, suggesting that those strategies sometimes benefit certain groups more than others. We report regression results for the remaining measures in Table 8 in Appendix E.4.
@@ 4.3 Model Comparison
Surprisingly, two of the largest models, i.e., Llama-3.3-70B and Gemma-3-27B, perform worst with respect to the Marked Personas framework and semantic diversity, while the smaller OLMo-2-7B, performs best (Fig. 4). 101010We observe the same pattern for classification accuracy (see Fig. 12 in Appendix E.3) and for the Social Media Bio task (see Fig. 9 in Appendix E.2).
@@ 5 Analysis of Marked Words
To complement the quantitative findings in Section 4, we also analyze the marked words identified in self-descriptions across all combinations of demographic group and prompt type. To this end, we compute the most frequent marked words per demographic group and present the top 10 in Table 2. In line with Cheng et al. (2023a), we observe that many of these words reflect patterns of markedness, essentialism, and othering in LLM responses. In particular, we identify four categories of marked words linked to problematic stereotypes as detailed in their work:
@@
Narrative of imposed resilience for Black personas (e.g., “resilience”, “strength”)
@@
Relationship to demographic identity as the primary frame of any non-White groups (e.g., “heritage”, “culture”)
@@
Conflation of Middle-Eastern identity with religiosity (e.g., “faith”, “muslim”)
@@
Disproportionate focus on gender identity for nonbinary personas (e.g., “gender”, “identity”)
For each category, we calculate the share of personas from the associated groups whose self-descriptions include such words and compare between prompt strategies. Figure 6 shows results for category 1 and 2: using names and/or the interview format consistently lowers the share of stereotyped persona descriptions across models.111111Results for additional stereotype categories follow the same pattern; full results and significance levels are provided in Appendix E.6.
@@ 6 Related Work
Personas in LLMs. Personas, widely used in Interaction Design to represent user archetypes Cooper et al. (2014), have recently been adapted to mold the behavior of language technologies like chatbots and LLMs. Personas in LLMs span e.g. psychometric, occupational, and demographic dimensions to guide model behavior and tailor responses to contexts or user needs Tseng et al. (2024). Inducing personas in LLMs effectively is an active area of research Jiang et al. (2023); Shu et al. (2024); Liu et al. (2024). We study demographic personas which can be used to create synthetic samples of particular human subpopulations Argyle et al. (2023).
@@ 7 Discussion and Conclusion
Our findings show that LLMs often stereotype marginalized groups, especially nonbinary, Hispanic, and Middle Eastern personas, highlighting ongoing challenges in representing demographic groups (Cheng et al., 2023b, a).
@@
Using last names for demographic priming also shows promise in mitigating biases and aligns with Wang et al. (2025), who found similar effects using first names. However, using first names to signal gender identity, can entrench biases towards certain groups, e.g, nonbinary people Gautam et al. (2024). We circumvent this by instead relying on titles (i.e, Mr., Ms., and Mx.) to signal gender. Nevertheless, we observe that models occasionally interpret last names in unintended ways (for example, some names associated with Black identity in the U.S. were linked to French residents), posing challenges to validity (Gautam et al., 2024). Thus, we suggest treating names as a potentially useful but ethically fraught tool, that requires careful implementation and critical reflection on its limitations.
A new observation is that LLMs switch language during simulation, but only when simulating Hispanic personas. To the best of our knowledge, this has not been studied before. Investigating whether this extends to other identities and languages, and how it relates to LLMs’ multilingual capabilities, is a promising direction for future work.
Finally, we stress the importance of clearly documenting and justifying choices around role adoption, demographic priming, and the specific descriptors used in persona prompts. To support future work, we release our code and datasets.121212https://github.com/dess-mannheim/prompt-makes-the-persona
@@
While our work lays a foundation for evaluating sociodemographic persona prompts, several limitations remain. First, the study is limited to English and focuses only on two commonly studied dimensions — gender and race (Sen et al., 2025) — which constrains generalizability to other demographic dimensions and languages. Additionally, because the persona prompts are constructed in a fully data-driven manner, the resulting set is not exhaustive and may miss other meaningful variations. This limitation also extends to the choice of demographic descriptors; for example, we do not systematically compare other demographic descriptors such as "girl" or "father", leaving such analysis for future work.
To assess the alignment with human responses on OpinionsQA, we prompt models using a multiple-choice format, reflecting the design of the original human survey. While this has become a standard evaluation paradigm in the field (Hwang et al., 2023; Moon et al., 2024), recent research has highlighted important limitations of this approach (Röttger et al., 2024), which also apply to our work. Additionally, we do not randomize or control the order of answer options due to computational constraints. This may influence model response patterns (Rupprecht et al., 2025) and, consequently, the observed alignment behavior. Lastly, while we considered the possibility of data contamination in OpinionsQA, we find it unlikely to be a confounding factor, as none of the models show particularly strong alignment on the task.
Finally, we always place the persona prompts in the user message, as our literature review found this to be the most common approach in prior studies. We anticipate that using the system prompt instead could reduce the frequency of role violations, particularly for Gemma-3-27b, which showed the highest rate of such violations (cf. Appendix E.1). A systematic comparison of both strategies is an interesting venue for future work.
@@
Names as Demographic Proxies.
Name-based prompting presents additional challenges. The racial categories and last names we use in this paper are based on U.S. Census data, biasing the study toward U.S.-centric naming conventions and limiting applicability to other geographic and cultural contexts. Moreover, focusing only on the most common last names risks capturing a narrow and potentially unrepresentative subset of a group. We further find that LLMs can interpret last names in unintended ways. For instance, last names associated with Black identity in the U.S. were occasionally interpreted as French residents (e.g., “Mr. Jeanbaptiste, a 35-year-old history enthusiast from Paris”), indicating that name-based priming may invoke unintended geographic or cultural associations. Crucially, we caution against relying on first names to infer or encode gender, as this risks misgendering individuals and excluding nonbinary and gender-diverse people; therefore, we use titles instead to signal gender. We refer readers to Gautam et al. (2024) for a comprehensive discussion of the limitations and ethical considerations involved in using personal names to infer or signal sociodemographic attributes.
@@ Acknowledgments
We thank Florian Lemmerich for the valuable input and helpful discussions in the early stages of this project. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG.
@@ Appendix B Sociodemographic Prompt Construction
Table 3 presents all prompt templates used in our experiments. Task-specific strings are indicated by a placeholder ([TASK]) and substituted as shown in Table 4. Placeholders indicating race/ethnicity through explicit descriptors ([R]) or last names ([NM]) are listed in Table 5. The remaining placeholders are populated with demographic descriptors as indicated in Tables 6 and 7.
Last Names.
We source last names for all racial and ethnic groups from Aher et al. (2023), except for Middle Eastern names, which are not included in that study. For Middle Eastern race/ethnicity, we instead select the most common last names from the three largest Middle Eastern countries, namely Turkey,131313https://forebears.io/turkey/surnames Iran,141414https://forebears.io/iran/surnames and Saudi-Arabia.151515https://forebears.io/saudi-arabia/surnames For each task, we include all 10 names per demographic group. In the open-ended tasks, each prompt is run with every individual name using 10 random seeds (instead of 100), resulting in 100 responses per demographic category.
All measures are computed on the combined set of responses for all names within the same demographic category (i.e., all names listed in a single row of Table 5). For semantic diversity and opinion distance, we additionally calculated the mean pairwise distance of embedding vectors and Wasserstein distance, respectively, for the responses associated with each individual name, and then averaged these values across names within the same demographic category. We found that this alternative approach produced results that were very similar to those obtained from the aggregated data, and, more importantly, the overall patterns remained consistent.
@@ Appendix C Computational Details
We use vLLM161616https://github.com/vllm-project/vllm version 0 (for the open-ended tasks) and version 1 (for OpinionsQA since it was faster) for LLM inference. We run the Llama-3.1-8B-Instruct and OLMo-2-1124-7B-Instruct models on a single Nvidia A100 (40GB) GPU with a total running time of approximately 9:50h (0:30h for self-descriptions, 0:20h for social media bios and 9:00h for OpinionsQA) per model. We run Llama-3.3-70B-Instruct, OLMo-2-0325-32B-Instruct and gemma-3-27b-it on two Nvidia H100 (80GB) GPUs in parallel. The total running time for generating all outputs with all three models is approximately 43:40h (5:40h for self-descriptions, 2h for social media bios, 36h for OpinionsQA).
For all models, we use the parameters recommended by the model creators. To generate answer option distributions for OpinionsQA, we extract temperature-scaled log probabilities, applying the temperature values recommended by the model creators.
@@ Appendix D Preprocessing of LLM Responses
For the open-text responses we apply a variety of measures to identify stereotypes and flattening effects, with the normative expectation that groups are not treated differently when generating self-descriptions or social media bios for them based on demographics alone. To accurately detect such patterns, we first remove all demographic markers, names, and pronouns from our prompts that have been repeated in the LLMs’ answer. Using a regex, we replace them with the string "[identity]", while ensuring that consecutive redactions are combined. For instance, we replace "I am Ms. Garcia, a …" with "I am [identity], a …", when "Ms." and "Garcia" were both part of our prompt.
@@ D.1 Identification of Role Violations
We further filter the set of free-text responses to exclude cases in which the LLMs failed to maintain their assigned role. This includes instances where the model refused to answer (e.g., "I cannot fulfill this request."), answered as an AI assistant (e.g., "Just a language model navigating the human world.") or partially reverted to an AI identity after initially adopting the persona (e.g., "I am a 32-year-old Black female. I identify as an AI…").
To identify such role violations, we used Qwen2.5-32B Bai et al. (2023)—a model not included in our main experiments to avoid self-preference bias in LLM annotations Wataoka et al. (2024)—for automated annotation. We then validated this automatic classification by conducting a small annotation study, with 300 randomly sampled LLM responses where three human annotators who are fluent in English verified the labels. Both, the human annotators and Qwen2.5-32B received a similar instruction with the grey part being only shown to the LLM: